OpenStudy (love_333):

Medal + fan The volumes of two similar solids are 729m^3 and 125m^3. The surface area of the larger solid is 324m^2. What is the surface area of the smaller solid? My Guess**A 56m^2 B 100m^2 C 500m^2 D 200m^2

OpenStudy (zachnovo15):

What do you think?

OpenStudy (love_333):

If I had to guess?... A?

OpenStudy (zachnovo15):

That's what it looks like the answer is.

OpenStudy (erak):

I kind of want to go with B, but just stick with A for now

OpenStudy (love_333):

Why were you thinking b? @Erak

OpenStudy (zachnovo15):

A kinda seems like the obvious answer

OpenStudy (zachnovo15):

I see where @Erak is coming from.

OpenStudy (erak):

Okay for the ratio between them because they are "similar solids", I will assume they are both cubes. cbrt(729) = 9 cbrt(125) = 5 The sides are in a 9:5 ratio, or change of 1.8 324 = 6s^2 s = sqrt(54) That's what I get for the sidelength, now to convert it to the smaller cube, divide by 1.8, so sqrt(54)/1.8 SA = 6s^2 SA = 6(sqrt(54)/1.8)^2 SA = 100m^2

OpenStudy (erak):

@jim_thompson5910 want to check my reasoning?

jimthompson5910 (jim_thompson5910):

The ratio of the volumes is 729:125 Take the cube root of both pieces to get 9:5 Notice how 9/5 = 1.8 So the length of the larger object is 1.8 times bigger than the corresponding length of the smaller object I.E. length of the larger object = 1.8*(length of the smaller object) ------------------------------------------------------------------ Square 1.8 to get 3.24 Let x = larger surface area y = smaller surface area The ratio of the surface areas is going to be (larger surface area)/(smaller surface area) = 3.24 x/y = 3.24 324/y = 3.24 324 = 3.24y y = 100 So you are correct. The smaller object has a surface area of 100 m^2

OpenStudy (erak):

@jim_thompson5910 thank you, i knew it couldn't be as simple as A